# Synergizing LLMs and Knowledge Graphs: A Novel Approach to Software Repository-Related Question Answering

### v1

@@ Abstract.
Software repositories contain valuable information for gaining insights into their development process. However, extracting insights from these repository data is time-consuming and requires technical expertise. While software engineering chatbots have been developed to facilitate natural language interactions with repositories, they struggle with understanding natural language and accurately retrieving relevant data. This study aims to improve the accuracy of LLM-based chatbots in answering repository-related questions by augmenting them with knowledge graphs. We achieve this in a two-step approach; (1) constructing a knowledge graph from the repository data and (2) synergizing the knowledge graph with LLM to allow for the natural language questions and answers. We curated a set of 20 questions with different complexities and evaluated our approach on five popular open-source projects. Our approach achieved an accuracy of 65%. We further investigated the limitations and identified six key issues, with the majority relating to the reasoning capability of the LLM. We experimented with a few-shot chain-of-thought prompting to determine if it could enhance our approach. This technique improved the overall accuracy to 84%. Our findings demonstrate the synergy between LLMs and knowledge graphs as a viable solution for making repository data accessible to both technical and non-technical stakeholders.
@@
ccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Software notations and tools
@@ 1. Introduction
Software repositories are rich sources of information essential to the software development process. This includes data on source code, documentation, issue tracking data, and commit histories (Vidoni, 2022). Analyzing this data can provide valuable insights about a project, such as developer activities and project evolution (Hassan, 2008). For instance, Begel and Zimmermann (2014) and Sharma et al. (2017) presented questions that software practitioners are interested in answering regarding their projects. Answering some of these questions requires mining and analyzing repository data. However, accessing and extracting meaningful insights from repositories is time-consuming and requires technical expertise (Abdellatif et al., 2020a; Banerjee and Cukic, 2015). For example, in a StackOverflow post (izkeros, 2022), a user seeking to calculate the number of lines changed since the last commit in a Git repository found that the solution required using specific Git commands like git diff --shortstat, which can be challenging for non-technical stakeholders. The technical knowledge and the time spent on such a task can be a barrier to software practitioners.
@@
Prior studies have attempted to address this challenge by developing software engineering chatbots that provide intuitive, natural language interfaces to software repositories (Abdellatif et al., 2020a; Abedu et al., 2024a). However, a key challenge in software engineering chatbot development lies in natural language understanding (NLU), as the chatbot should accurately interpret user questions and map them to appropriate data retrieval actions (Abdellatif et al., 2020b). Additionally, the NLU approach to chatbot development fails when the NLU model is not trained on the intent of the user’s question. Each intent is typically mapped to a predefined action within the chatbot’s framework (Ramesh et al., 2017; Adamopoulou and Moussiades, 2020). However, it is often impractical to define actions for every possible intent, especially as user requirements evolve, limiting the chatbot’s functionality and adaptability. Large language models (LLMs) have demonstrated remarkable capabilities in understanding natural language and identifying the intents of input texts (Shah et al., 2024). Nonetheless, leveraging LLMs to build chatbots for repository question answering using the retrieval augmented generation (RAG) approach has proved challenging. Abedu et al. (2024a) reported that LLM-based RAG chatbots fail to retrieve accurate data to answer repository-related questions 83.3% of the time.
@@
Knowledge graphs have the potential to enhance LLMs with external data to generate contextually relevant responses (Pan et al., 2024; Abu-Rasheed et al., 2024). Knowledge graphs are structured representations that model entities and their relationships, enabling enhanced semantic and structural understanding and reasoning (Ji et al., 2022; Chen et al., 2020; Xie et al., 2016). Due to the structured nature of software repositories, prior studies have modeled software repositories as knowledge graphs to solve various problems in the software engineering domain (Malik et al., 2024; Zhao et al., 2019; Ma et al., 2024). Motivated by this, we aim to improve the accuracy of LLM-based chatbots in answering repository-related questions by synergizing LLMs with knowledge graphs.
@@
Our approach relies on the ability of LLMs to generate correct graph queries from natural language. Therefore, we conducted an exploratory analysis to assess and identify the most efficient state-of-the-art LLM for graph query generation from natural language input. We evaluated several LLMs, including GPT-4o (OpenAI, 2024b), Llama3 (Meta, 2024), and Claude3.5 (Anthropic, 2024), on their ability to generate accurate graph queries which we operationalize using Cypher (Francis et al., 2018) from natural language text and found GPT-4o the most accurate in our context. We integrated GPT-4o with our knowledge graph-based approach and evaluated the approach on five popular open-source repositories (namely, AutoGPT, Bootstrap, Ohmyzsh, React, Vue) using 20 questions curated from Abdellatif et al. (2020a). To further evaluate the performance of our approach on the complexity of the questions, we grouped the questions into three difficulty levels based on the number of relationships in the knowledge graph required to retrieve the relevant data. Specifically, our evaluation aims to answer the following research questions:
@@
  • We provide empirical evidence demonstrating the capabilities of LLMs in generating Cypher queries for querying knowledge graphs.
  • We discuss the limitations of augmenting LLMs with knowledge graphs.
@@ 2.1. Software Repositories
2.2. Knowledge Graphs
@@
2.3. Software Engineering Chatbots
Chatbots are conversational assistants designed to assist with specific tasks by interacting with users through natural language (Ramesh et al., 2017). They aim to facilitate access to information, automate routine tasks, and support collaboration among team members (Abdellatif et al., 2022). Chatbots are increasingly becoming popular in the software engineering domain to accomplish specific software engineering tasks. For instance, Abdellatif et al. (2020a) proposed MSRBot, using a bot layered on top of software repositories to automate and simplify the extraction of useful information from the repository. Bradley et al. (2018) proposed Devy, a Conversational Developer Assistant that enables developers to focus on high-level tasks by reducing the need for manual low-level commands across various tools. Dominic et al. (2020) proposed a conversational bot to support newcomers in onboarding to open-source projects by recommending suitable projects, resources, and mentors. Okanović et al. (2020) proposed PerformoBot, a chatbot that guides developers through configuring and executing load tests via natural language conversations. Also, Abedu et al. (2024a) developed an LLM-based chatbot to answer questions related to software repositories. Their LLM-based chatbot, which used the RAG approach, failed to retrieve the relevant data needed to answer questions in their evaluation questions most of the time.
The increasing application of chatbots in software engineering and LLMs in chatbots like ChatGPT and BARD motivates our work to improve the accuracy of LLMs in software engineering chatbots. To the best of our knowledge, this is the first software repository question-answering approach based on knowledge graphs and LLMs. Abdellatif et al. (2020a) and Abedu et al. (2024a) are the closest to our work. However, the approach by Abdellatif et al. (2020a) supports a limited number of intents. LLMs have demonstrated the ability to identify intents of input text (Shah et al., 2024), using LLMs in this study extends the impact of Abdellatif et al. (2020a) by supporting more user questions with variable intents. Our work also seeks to address the challenge reported by Abedu et al. (2024a) regarding LLM-based chatbots inaccurately retrieving relevant data as context by synergizing the LLM with knowledge graphs which has the potential to enhance LLMs with relevant external data (Pan et al., 2024; Abu-Rasheed et al., 2024).
@@ 3.1. Knowledge Graph Constructor
The Knowledge Graph Constructor component aims to connect the entities in the software repository to form the repository knowledge graph. Using a knowledge graph allows us to model the complex relationships between the repository entities, facilitating analysis and inference of the repository data. The knowledge graph for a GitHub repository is complex considering the official GitHub schema (GitHub, 2024c). As a result, in this study, we limit the knowledge graph to four entities (namely, Users, Commits, Issues, and Files) and their relationships to be able to capture the nuance of the repository metadata without too much complexity. The knowledge graph constructor collects the following types of data as mentioned:
@@
(Commit, Parent_of, Commit) Indicates that a Commit is the parent of another Commit.
@@ 4.2. Evaluation Questions
In evaluating the approach, we curated the questions by Abdellatif et al. (2020a), which they collected from 12 users interacting with software repositories to access various information for the completion of tasks assigned to them. These tasks include finding answers to questions that are commonly asked by developers and non-technical stakeholders, for instance, finding the commit that introduced a bug or the developer that fixed the most bugs (Sharma et al., 2017; Begel and Zimmermann, 2014). The users asked 165 questions representing 10 distinct intents, where each intent refers to the mapping between the user’s question and a predefined action to be taken to complete the task (Adamopoulou and Moussiades, 2020). Executing all 165 questions would be expensive; therefore, we limited our selection to two questions per intent, resulting in a final evaluation set of 20 questions. The two questions were selected for each intent based on the clarity of their phrasing. This approach reduces the cost of executions while ensuring variety by covering each of the 10 intents with two variations of questions.
@@ 4.3. Implementation
We implement the approach discussed in Section 3 using the Python and Langchain framework. The Knowledge Graph Constructor begins the process by collecting data from the software repository using the GitHub GraphQL API (GitHub, 2024a). We opted for the GraphQL API over the REST API (GitHub, 2024b) because GraphQL allows us to specify precisely the data we need in a single request, reducing the noise and the size of the document returned compared to the REST API and improving processing efficiency (Brito et al., 2019). The collected data included information on users, commits, issues, and files (see Section 3.1). After collecting the data, we implemented the relationship between the entities following the schema defined in Figure 2. To store and manage the knowledge graph, we utilize the Neo4j database, which is known for its robustness and maturity in handling graph data structures and widely adopted in prior studies (Malik et al., 2024). Its compatibility with the Cypher query language enables efficient querying and manipulation of the graph data (Francis et al., 2018).

### v2

@@ Abstract.
Software repositories contain valuable information for understanding the development process. However, extracting insights from repository data is time-consuming and requires technical expertise. While software engineering chatbots support natural language interactions with repositories, chatbots struggle to understand questions beyond their trained intents and to accurately retrieve the relevant data. This study aims to improve the accuracy of LLM-based chatbots in answering repository-related questions by augmenting them with knowledge graphs. We use a two-step approach: constructing a knowledge graph from repository data, and synergizing the knowledge graph with an LLM to handle natural language questions and answers. We curated 150 questions of varying complexity and evaluated the approach on five popular open-source projects. Our initial results revealed the limitations of the approach, with most errors due to the reasoning ability of the LLM. We therefore applied few-shot chain-of-thought prompting, which improved accuracy to 84%. We also compared against baselines (MSRBot and GPT-4o-search-preview), and our approach performed significantly better. In a task-based user study with 20 participants, users completed more tasks correctly and in less time with our approach, and they reported that it was useful. Our findings demonstrate that LLMs and knowledge graphs are a viable solution for making repository data accessible.
@@
$c$s: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Software notations and tools
@@ 1. Introduction
Software repositories are rich sources of information essential to the software development process. This includes data on source code, documentation, issue tracking data, and commit histories (Vidoni, 2022). Analyzing this data can provide valuable insights about a project, such as developer activities and project evolution (Hassan, 2008). For instance, Begel and Zimmermann (2014) and Sharma et al. (2017) presented questions that software practitioners are interested in answering regarding their projects. Question answering is a core part of day-to-day development: prior studies show that developers continuously seek answers about project state, processes, and code to make progress, underscoring the need for tools that surface precise, timely answers (Begel and Zimmermann, 2014; Ko et al., 2007). Answering some of these questions requires mining and analyzing repository data. However, accessing and extracting meaningful insights from repositories is time-consuming and requires technical expertise (Abdellatif et al., 2020a; Banerjee and Cukic, 2015). For example, in a StackOverflow post (izkeros, 2022), a user seeking to calculate the number of lines changed since the last commit in a Git repository found that the solution required using specific Git commands like git diff --shortstat, which can be challenging for non-technical stakeholders. The technical knowledge and the time spent on such a task can be a barrier to software practitioners.
@@
Prior studies have attempted to address this challenge by developing software engineering chatbots that provide intuitive, natural language interfaces to software repositories (Abdellatif et al., 2020a; Abedu et al., 2024a). However, a key challenge in software engineering chatbot development lies in natural language understanding (NLU), as the chatbot should accurately interpret user questions and map them to appropriate data retrieval actions (Abdellatif et al., 2020b). Additionally, the NLU approach to chatbot development fails when the NLU model is not trained on the intent of the user’s question. Each intent is typically mapped to a predefined action within the chatbot’s framework (Ramesh et al., 2017; Adamopoulou and Moussiades, 2020). However, it is often impractical to define actions for every possible intent, especially as user requirements evolve, limiting the chatbot’s functionality and adaptability. Large language models (LLMs) have demonstrated remarkable capabilities in understanding natural language and identifying the intents of input texts (Shah et al., 2024). Nonetheless, leveraging LLMs to build chatbots for repository question answering using the naive retrieval-augmented generation (RAG) approach has proved challenging. Abedu et al. (2024a) reported that LLM-based RAG chatbots failed to retrieve accurate data to answer repository-related questions 83.3% of the time.
@@
Knowledge graphs have the potential to enhance LLMs with external data to generate contextually relevant responses (Pan et al., 2024; Abu-Rasheed et al., 2024). Knowledge graphs are structured representations that model entities and their relationships, enabling enhanced semantic and structural understanding and reasoning (Ji et al., 2022; Chen et al., 2020; Xie et al., 2016). Due to the structured nature of software repositories, prior studies have modelled software repositories as knowledge graphs to solve various problems in the software engineering domain (Malik et al., 2024; Zhao et al., 2019; Ma et al., 2024). Motivated by this, we aim to improve the accuracy of LLM-based chatbots in answering repository-related questions by synergizing LLMs with knowledge graphs. In this study, similar to prior studies, we limit the knowledge graph modelling to Git metadata (Malik et al., 2024) and focus on answering repository-related questions that are limited to Git metadata, similar to (Abdellatif et al., 2020a; Abedu et al., 2024a).
@@
Our approach relies on the ability of LLMs to generate correct graph queries from natural language. Therefore, we conducted an exploratory analysis to assess and identify the most efficient state-of-the-art LLM for graph query generation from natural language input. We evaluated several LLMs, including GPT-4o (OpenAI, 2024b), Llama3 (Meta, 2024), and Claude3.5 (Anthropic, 2024), on their ability to generate accurate graph queries which we operationalize using Cypher (Francis et al., 2018) from natural language text and found GPT-4o the most accurate in our context. We integrated GPT-4o with the knowledge graph–based approach and evaluated it on five popular open-source repositories (AutoGPT, Bootstrap, Ohmyzsh, React, Vue) using 150 questions curated from Abdellatif et al. (2020a). We first selected a subset of 20 questions spanning the intents and difficulty levels to run the initial evaluation and identify failure modes; we then ran the best configuration on the full set of 150 questions. Questions were grouped into three difficulty levels based on the number of relationships in the knowledge graph needed to retrieve the data. Each question was executed five times with a majority vote for correctness. We compared against MSRBot and GPT-4o-search-preview as baselines and conducted a task-based user study with 20 participants. Specifically, our evaluation aims to answer the following research questions:
@@
  • RQ4:
    How do users perceive the usefulness of our approach in assisting them to answer repository-related questions? We find that the study participants completed more tasks correctly and in less time with the chatbot than with their usual methods, and they perceived the chatbot as useful and time-saving.
@@
  • We provide empirical evidence demonstrating the capabilities of LLMs in generating Cypher queries for querying software repository data stored/represented in knowledge graphs.
  • We discuss the limitations of augmenting LLMs with software repository data stored/represented in knowledge graphs.
@@ 2.1. Software Repositories
2.2. Software Engineering Chatbots
Chatbots are conversational assistants designed to assist with specific tasks by interacting with users through natural language (Ramesh et al., 2017). They aim to facilitate access to information, automate routine tasks, and support collaboration among team members (Abdellatif et al., 2022). Evidence from recent studies underscores that question-answering has become central to developer workflows: controlled and field studies report sizable productivity gains from conversational assistants, while surveys show widespread but cautious adoption of AI assistants for code search, explanation, and guidance (Khojah et al., 2024; Peng et al., 2023; Stackoverflow, 2024). Chatbots are increasingly becoming popular in the software engineering domain to accomplish specific software engineering tasks. For instance, Abdellatif et al. (2020a) proposed MSRBot, using a bot layered on top of software repositories to automate and simplify the extraction of useful information from the repository. Bradley et al. (2018) proposed Devy, a Conversational Developer Assistant that enables developers to focus on high-level tasks by reducing the need for manual low-level commands across various tools. Dominic et al. (2020) proposed a conversational bot to support newcomers in onboarding to open-source projects by recommending suitable projects, resources, and mentors. Okanović et al. (2020) proposed PerformoBot, a chatbot that guides developers through configuring and executing load tests via natural language conversations. Also, Abedu et al. (2024a) developed an LLM-based chatbot to answer questions related to software repositories. Their LLM-based chatbot, which used the RAG approach, failed to retrieve the relevant data needed to answer questions in their evaluation questions most of the time. Beyond these task-specific bots, recent code-aware assistants such as GitHub Copilot Chat, CodeT5+, and StarCoder bring conversational capabilities into the IDE, supporting code generation, explanation, and navigation; however, they primarily target source-level reasoning rather than repository metadata Q&A across issues and commits (Zhuo et al., 2025; Wang et al., 2023a).
The increasing application of chatbots in software engineering and LLMs in chatbots like ChatGPT and BARD motivates our work to improve the accuracy of LLMs in software engineering chatbots. To the best of our knowledge, this is the first software repository question-answering approach based on knowledge graphs and LLMs.
Abdellatif et al. (2020a) and Abedu et al. (2024a) are closest to our work. Like MSRBot (Abdellatif et al., 2020a), we follow a similar interaction pattern, where a user asks a question; the approach queries a repository’s data and then generates an answer for the user. Also, we evaluate our approach using the evaluation dataset from Abdellatif et al. (2020a). However, unlike Abdellatif et al. (2020a)’s fixed intent/entity pipeline, our work uses an LLM generator with a repository knowledge graph, enabling support for more user questions with variable intents. Abedu et al. (2024a) showed that RAG-based LLM chatbots often fail because they retrieve irrelevant context; we address this by grounding the LLM in a structured knowledge graph that supplies precise repository facts rather than unstructured snippets, reducing retrieval mismatch and improving answer reliability (Pan et al., 2024; Abu-Rasheed et al., 2024). While Copilot-style systems allow users to interact with their repository at the code level, our focus is on allowing users to interact at the repository-metadata level.
2.3. Knowledge Graphs and Large Language Models
@@
LLMs are effective at natural language but can hallucinate, lack provenance, and provide out-of-date information. Knowledge graphs provide explicit and verifiable facts with typed relations and support multi-hop querying. Integrating the two technologies improves LLM outputs and accuracy. Prior studies on this combination outline these benefits. For instance, Li et al. (2024) converts tables into graphs and guides step-by-step reasoning, which filters noise and yields higher QA accuracy than text-based baselines. Xu et al. (2025) retrieve entities, relations, and subgraphs and align them before prompting, which improves answer accuracy and logical form over text retrieval. Lavrinovics et al. (2025) review knowledge graph-based strategies that reduce hallucinations during pretraining, inference, and post-generation, and they call for stronger evaluation. Hogan et al. (2025) show that systems that delegate multi-hop facts to knowledge graphs, freshness to search, and fluency to LLMs deliver more reliable answers. Sequeda et al. (2025) argue that knowledge graphs enable trust and provenance in enterprise QA, and they report better auditability and governance when answers are checked against knowledge graph facts.
For repository question answering, combining LLMs with a knowledge graph is a viable approach because a repository knowledge graph provides verifiable facts that ground the LLMs’ responses and improve answer accuracy. To the best of our knowledge, this is the first study to address repository question answering using an LLM enhanced with a knowledge graph.
2.4. Prompt Engineering for LLMs in Software Engineering
Large language models are widely used for software engineering tasks such as code generation, summarization, and commenting. Prompt engineering guides these models with task-specific instructions without changing model weights. A recent systematic review by Hou et al. (2024) lists the common prompt engineering techniques used in recent software engineering studies. These include zero-shot, few-shot, chain-of-thought (CoT), automatic prompt engineering (APE), prompt-based continuous prompting (PromptCS), Chain-of-Code (CoC), modular-of-thought (MoT), and structured chain-of-thought (SCoT), and reports the broad use of few-shot and CoT prompting (Hou et al., 2024).
Geng et al. (2024) investigate multi-intent comment generation and report that providing ten or more in-context examples in the prompt (few-shot learning) enables diverse comments and outperforms a supervised baseline. Xu et al. (2024) propose UniLog for automatic logging and show that five demonstration examples in the prompt (few-shot learning), with example selection and ordering, improve insertion quality without model tuning. Wu et al. (2023) study security repair and use prompts that mark buggy and fixed regions (e.g., “BUG:”/“FIXED:”), finding that success is limited and mostly on simple cases.
Shin et al. (2025) compares GPT-4 under basic, in-context, and task-specific prompts to fine-tuned models across summarization, generation, and translation, and reports that prompt-engineered GPT-4 is on par with fine-tuned models and offers ease of use, especially with a natural language interface. These studies show that prompt engineering can improve results of LLM applications on SE tasks, especially with few-shot and CoT prompting (Hou et al., 2024; Shin et al., 2025).
@@ 3.1. Knowledge Graph Constructor
The Knowledge Graph Constructor component aims to connect the entities in the software repository to form the repository knowledge graph. Using a knowledge graph allows us to model the complex relationships between the repository entities, facilitating analysis and inference of the repository data. Given the size of the official GitHub schema (GitHub, 2024c), we restrict the graph to four entities (Users, Commits, Issues, and Files) and their relationships. These four entities capture the core workflows we evaluate (developer activity, commit history, and defect lifecycle), enabling multi-hop analysis while keeping the schema manageable for verification and traceability of responses. The knowledge graph constructor collects the following types of data:
@@
(Commit, Parentof, Commit) Indicates that a Commit is the parent of another Commit.
@@ 4.2. Evaluation Questions
In evaluating the approach, we curated the questions by Abdellatif et al. (2020a), which they collected from 12 users interacting with software repositories to access various information for the completion of tasks assigned to them. These tasks include finding answers to questions that are commonly asked by developers and non-technical stakeholders, for instance, finding the commit that introduced a bug or the developer that fixed the most bugs (Sharma et al., 2017; Begel and Zimmermann, 2014). The users asked 165 questions representing 10 distinct intents, where each intent refers to the mapping between the user’s question and a predefined action to be taken to complete the task (Adamopoulou and Moussiades, 2020).
Out of the 165 questions in the dataset, we exclude 15 questions that are not relevant to the evaluation. For example, “What’s your name?”. We use the remaining questions as templates for forming the evaluation questions for each project. This involves inserting project-specific parameters into the question template to form the question. For instance, in the original dataset, for the question “Who fixed the most bugs in the file HibernateEntityManager?”, we replace the file HibernateEntityManager with a file specific to the project we are evaluating. Executing all 150 questions for each project would be expensive; therefore, we create a subset of the dataset by selecting two questions per intent, resulting in an evaluation set of 20 question templates. The two questions were selected for each intent based on the clarity of their phrasing. This approach reduces the cost of executions while ensuring variety by covering each of the 10 intents with two variations of questions.
@@ 4.3. Implementation
We implement the approach discussed in Section 3 using Python and the Langchain framework. The Knowledge Graph Constructor begins the process by collecting data from the software repository using the GitHub GraphQL API (GitHub, 2024a). We opted for the GraphQL API over the REST API (GitHub, 2024b) because GraphQL allows us to specify precisely the data we need in a single request, reducing the noise and the size of the document returned compared to the REST API and improving processing efficiency (Brito et al., 2019). The collected data included information on users, commits, issues, and files (see Section 3.1). After collecting the data, we implemented the relationship between the entities following the schema defined in Figure 2. To store and manage the knowledge graph, we utilize the Neo4j database, which is known for its robustness and maturity in handling graph data structures and widely adopted in prior studies (Malik et al., 2024). Its compatibility with the Cypher query language enables efficient querying and manipulation of the graph data (Francis et al., 2018).
@@
4.4. Baseline Set Up
We benchmark our approach against two baseline approaches, the MSRBot framework (Abdellatif et al., 2020a) and the GPT-4o model with web searching capability. The MSRBot framework is an intent-based chatbot framework that maps user questions to predefined intents and also extracts the relevant entities from the question to retrieve the relevant repository data. In the implementation of MSRBot, Google Dialogflow was used for natural language understanding. However, the authors explained that other NLU platforms can be used in the framework and further evaluated different NLU platforms (Abdellatif et al., 2022). In the study, Abdellatif et al. (2022) found that Rasa NLU produces stable confidence scores (median ¿ 0.91) and generally outperforms Dialogflow in confidence reliability. As a result, in this study we implement the MSRBot framework with the Rasa NLU.
The second baseline approach we consider is an open-domain LLM-based chatbot that can handle repository-related questions by augmenting a large language model with retrieval. A common method for this is RAG. The RAG approach integrates an LLM with a knowledge base: user queries trigger a search in an indexed repository dataset, and the retrieved documents are then used by the LLM to generate an answer. In principle, RAG allows the bot to provide up-to-date, factual answers even about dynamic repository data. We initially considered a RAG-based baseline, similar to the one explored by Abedu et al. (2024a) for mining software repositories. However, we do not include the RAG baseline in our evaluation. The reason is that the RAG chatbot showed poor accuracy, failing to answer most of the questions in the MSRBot dataset. The authors explained the low accuracy as a result of retrieving irrelevant data or failing to generate correct answers. This outcome highlights that a naive RAG implementation is ineffective for this evaluation, as reported by Abedu et al. (2024a). Given this finding, a RAG baseline would not be a competitive baseline. Instead, we replace the RAG approach with an LLM with web-search capability baseline: GPT-4o with web search capabilities (the GPT-4o-search-preview model). This baseline uses OpenAI’s GPT-4o, which is a state-of-the-art LLM, augmented with the ability to perform live web searches. The GPT-4 model can thus retrieve current information from online sources (including public GitHub repositories) and incorporate that knowledge into its responses. This serves a similar purpose to the RAG approach, providing the language model with access to relevant up-to-date information.
4.5. Survey Design
To evaluate the usefulness of our approach in assisting users with repository-related questions, we designed a task-based user study. This study follows a methodology similar to that of Abdellatif et al. (2020a). Participants are asked to perform a set of software repository-related tasks: first, using conventional tools (baseline), and then using our chatbot-based approach. By comparing the performance of the tasks in these two settings, we assess the usefulness of our approach for task accuracy and efficiency.
We adopted the 10 tasks performed by participants in Abdellatif et al. (2020a). These tasks cover a range of questions developers ask about the repositories. To avoid overburdening participants, we divided the ten tasks into two sets of five. We randomly selected two projects (Vue and Ohmyzsh) from our study projects discussed in Section 4.1 as the subjects for these tasks. Participants were then randomly split into two groups (Group A and Group B). Group A was assigned the first set of five tasks on the Vue project, and Group B was assigned the second set of five tasks on the Ohmyzsh project. The groupings helped ensure that no participant had to perform all ten tasks and that each participant only dealt with one project’s context, reducing fatigue and learning effects.
We developed and deployed a web-based chatbot application [^1] https://repochattool.streamlit.app/ implementing our approach, which participants used to interact with the projects. The study was administered via an online survey (Qualtrics), structured into four sections. The first section collects demographic information about the participants. This offers us insight into the professional background, the number of years of experience the participants have using Git tools, and how often they use the Git tools.
In the second section, participants were presented with five tasks and instructed to complete each task using any tools or methods of their choice (except our chatbot). They could use Git command-line queries, repository browsing, writing scripts, web searches, or AI assistants, as desired. In order for participants not to spend all their effort on a single task and keep within the allotted time of 60 minutes for the survey, we asked the participants to spend not more than 6 minutes per task (i.e., approximately 30 minutes total for the five tasks in this section). After completing each task, the participants provided their answer (the outcome of the task) and a brief description of how they completed the task. The description allowed us to understand the effort and techniques adopted by the participants to complete the tasks. We did not impose strict tool restrictions in order to emulate a realistic scenario where developers can leverage all resources at their disposal. This baseline phase established a point of comparison for task difficulty and time taken without the assistance of our tool.
In the third section, participants performed the same set of five tasks, but this time using our chatbot. We asked the participants to phrase each task as a question in their own words to the chatbot. For each task, the survey asked them to copy the exact question they posed to the chatbot and the answer the chatbot returned. After receiving the chatbot’s answer, participants were shown the expected correct answer (ground truth) for that task and asked to evaluate the chatbot’s response. They indicated whether the bot’s answer was Correct, Partially Correct, or Incorrect, and could optionally provide comments explaining their choice. We captured the time taken to complete each task with the chatbot as well. This section allowed us to assess the chatbot’s effectiveness in answering the questions directly compared to the baseline.
In the final section, participants answered a brief questionnaire about their overall experience with the chatbot. They rated, on a 5-point Likert scale, the chatbot’s accuracy, usefulness, time-saving capability, and their overall satisfaction with the chatbot (1 = Strongly Disagree to 5 = Strongly Agree for each statement). They were also given an open-ended text box to provide any additional feedback, comments, or suggestions about the chatbot. This feedback provides insights into the user-perceived benefits or issues with our approach.
The survey was configured to automatically record the time spent on each task (using Qualtrics’ timing features), giving us precise measurements for task completion times in both the baseline and chatbot settings. We piloted the survey with two participants. We used the feedback from these participants to improve and refine the survey for clarity before administering it to the other participants. For the distribution of the survey, we reached out to our network of contacts from both academia and industry to participate. We also adopted a snowball technique by asking the participants to share with other professionals with Git experience in their network.